# Deep generative modelling of canonical ensemble with differentiable thermal properties

### v1

@@ C.2 Estimation of the heat capacity using first-order differentiation
where $q_{\theta}(\mathbf{x},\beta)$ is the distributino of the VaTD model. When the loss is low enough, we assume the distribution of VaTD has the form of the Boltzmann distribution, i.e.,
@@
where $E_{\theta}$ is the energy function learned by the VaTD. Then, Eq. (S32) can be reformulated into
@@
From Eq. (4), when the model is trained to be a Boltzmann distribution, $\log\bar{\mathcal{Z}}$ is a constant for any $\bm{x}$. In that case, the former term in Eq. (S34) disappears, and the latter is related to the heat capacity, so the heat capacity can be written as
$C_{v}=\beta\frac{\partial}{\partial\beta}\left[\mathop{\mathbb{E}}_{\mathbf{x}\sim q_{\theta}(\cdot,\beta)}\log q_{\theta}(\mathbf{x},\beta)\right]\approx\beta^{2}(\langle E^{2}_{\theta}\rangle-\langle E_{\theta}\rangle^{2}).$ (S35)
Appendix D Supplementary for the 2D Ising Model Experiment
D.1 Architecture and hyperparameters
The architecture we used is a standard ResNet [25] structure with the CNN layer replaced by the masked CNN layer of [24]. For training, we used the Adam optimizer [42], whose learning rate was reduced periodically. Because the reinforce estimation tends to have high variance, we clipped the gradient values that are greater than a maximum value. All structural and training parameters can be found in Tab. S1. Except for the configuration plots, all the results showed in the Ising model simulation in the main text are estimated using a batch size of $7000$.
Table S1: Hyperparameters of the PixelCNN model used in the 2D Ising model experiment.
@@ D.1 Architecture and hyperparameters
Activation function RELU [44]
@@
D.2 Statistical averages with direct sampling
In this section, we compared the numerical results from the differentiation of the free energy with those from the statistical averaging with direct sampling. As the model in Fig. 1 of the main text has a too-low loss value, one may think the condition of Eq. (S30) is nearly satisfied, so a clear difference is not visible. To better illustrate it numerically, we used a saved model near the end of the training which has a slightly higher loss value.
In Fig. S1, using this secondary model, we plotted the heat capacity values of the 2D Ising model using automatic differentiation and statisitcal averaging with direct sampling. Compared with the automatic differentiation results, statistical averages show a larger deviation. In this plot, a batch of $7000$ samples was generated by the model at each temperature point.
Figure S1: The estimated heat capacity of the 2D Ising model on a $16\times 16$ PBC square lattice. The estimation results using differentiation of free energy (labeled as autoDiff) and statistical averaging with direct sampling (labeled as statistical) are compared with the MCMC results.
Appendix E Supplementary for the 2D XY Model Experiment
E.1 Architecture and hyperparameters
Our architecture consists of multiple layers of the piece-wise cubic spline transformation flow [35]. The cubic spline transformation uses interpolation points to create a monotonically increasing function to parameterize the invertible transformation. The coordinates and gradients of these interpolation points were computed using a standard ResNet [25]. The cubic spline flow also needs a way to bisect the variables. As the 2D XY configuration is a 2D array, we used the checkerboard pattern to separate it into two parts with an equal amount of variables. For training, we used the Adam optimizer [42], and we decreased the learning rate periodically. All structural and training parameters can be found in Tab. S2. Except for the configuration plots, all the results showed in the XY model experiment of the main text were estimated using a batch size of $1000$.
Table S2: Hyperparameters of the NF model used in the 2D XY model experiment.
@@ E.1 Architecture and hyperparameters
Figure S2: The estimated heat capacity of the 2D XY model on a $16\times 16$ PBC square lattice. The estimation results using differentiation of free energy (labeled as autoDiff) and statistical averaging with direct sampling (labeled as statistical) are compared with the MCMC results.
E.2 Statistical averages with direct sampling
Using this XY case, we further demonstrated numerically the difference between automatic differentiation and direct sampling. In Fig. S2, we gave the statistical averaging results of the heat capacity compared with differentiation results. In this plot, a batch of $1000$ samples was directly sampled by the model at each temperature point, for both the statistical averages and differentiation results. One can see that the statistical averages have a larger variance using the same number of samples.

### v2

@@ D.4 Estimation of the heat capacity using first-order derivative
where $q_{\theta}(\mathbf{x},\beta)$ is the distribution of the VaTD model. When the loss is low and STD is small, we assume the distribution of VaTD model has the form of the Boltzmann distribution, i.e.,
@@
where $E_{\theta}$ is the energy function learned by the VaTD model, and because of a small STD, $\bar{\mathcal{Z}}$ is a stable estimate of the partition function which doesn’t depend on $\bm{x}$. Then, Eq. (S33) can be reformulated into
@@
As we consider $\log\bar{\mathcal{Z}}$ is a constant for any $\bm{x}$, the former term in Eq. (S35) disappears, and the latter is related to the heat capacity of the learned energy function. So, the heat capacity estimate can be written as
$C_{v}=\beta\frac{\partial}{\partial\beta}\left[\mathop{\mathbb{E}}_{\mathbf{x}\sim q_{\theta}(\cdot,\beta)}\log q_{\theta}(\mathbf{x},\beta)\right]\approx\beta^{2}(\langle E^{2}_{\theta}\rangle-\langle E_{\theta}\rangle^{2}),$ (S36)
which can be estimated using only the first-order derivative.
Appendix E Supplementary for the 2D Ising Model Experiment
E.1 Architecture and hyperparameters
The architecture we used is a standard ResNet [resnet] structure with the CNN layer replaced by the masked CNN layer of Ref. [pixelRCNN]. For training, we used the Adam optimizer [adam], whose learning rate was reduced periodically. Because the REINFORCE estimation tends to have high variance, we clipped the gradient values that are greater than a maximum value. All structural and training parameters can be found in Tab. S2.
We used a collection of 8 temperature points at $\beta=0.5,0.8,0.9,1,1.1,1.2,1.3,1.8$ as the evaluation loss. We first trained the model till this evaluation loss reached equilibrium. From those saved models at equilibrium, we then selected the optimal model with the smallest maximum difference with the exact results over the temperature range of $[0.1,1.0]$
For the plots in the main text, we used Eq. (S32) to control the convergence error to a level of $5\times 10^{-3}$.
Table S2: Hyperparameters of the VaTD PixelCNN model used in the 2D Ising model experiment.
@@ E.1 Architecture and hyperparameters
Activation function RELU [relu]
@@
E.2 Statistical averages with direct sampling
In this section, we compared the numerical results from the free energy derivative estimation method with those from the statistical averaging with direct sampling. As the model used to plot Fig. 1 of the main text has a very low loss value, one may think the condition of Eq. (S30) is nearly satisfied, so a clear difference between these two methods is not visible. To better illustrate it numerically, we used a saved model near the end of the training which has a slightly higher loss value.
In Fig. S4, using this secondary model, we plotted the heat capacity values of the 2D Ising model using free energy derivative estimation and statistical averaging with direct sampling. Compared with the derivative estimation results, statistical averages show a larger deviation. In this plot, for both the derivative estimation and statistical averaging, a batch of $7000$ samples was generated by the model at each temperature point.
Figure S4: The estimated heat capacity of the 2D Ising model on a $16\times 16$ square lattice with PBC. The estimation results using the differentiation of free energy (labeled as autoDiff) and statistical averaging with direct sampling (labeled as statistical) are compared with the MCMC results.
Appendix F Supplementary for the 2D XY Model Experiment
F.1 Architecture and hyperparameters
Our architecture consists of multiple coupling layers of the piece-wise cubic spline transformation [cubic]. The cubic spline transformation uses interpolation points to create a monotonically increasing function to parameterize an invertible transformation. The coordinates and gradients of these interpolation points were computed using a standard ResNet [resnet]. This cubic spline flow also needs a way to bisect the variables. As the 2D XY configuration is a 2D array, we used the checkerboard pattern to separate it into two parts with an equal amount of variables. For training, we used the Adam optimizer [adam], and we decreased the learning rate periodically. All structural and training parameters can be found in Tab. S3.
We used a collection of 8 temperature points at $\beta=0.5,0.75,0.85,0.9,1.,1.1,1.25,2.4$ as the evaluation loss. We first trained the model till this evaluation loss reached equilibrium. From these saved models at equilibrium, we then selected the optimal model with the smallest maximum STD over the temperature range of $[0.5,1.5]$.
For the plots in the main text, we used Eq. (S32) to control the convergence error to a level of $5\times 10^{-3}$.
Table S3: Hyperparameters of the VaTD NF model used in the 2D XY model experiment.
@@ F.1 Architecture and hyperparameters
Figure S5: The estimated heat capacity of the 2D XY model on a $16\times 16$ square lattice with PBC. The estimation results using the differentiation of free energy (labeled as autoDiff) and statistical averaging with direct sampling (labeled as statistical) are compared with the MCMC results.
F.2 Statistical averages with direct sampling
Using this XY case, we further demonstrated the difference between the free energy derivative estimation method and statistical averaging with direct sampling. In Fig. S5, we gave the statistical averaging results of the heat capacity compared with derivative estimation results. In this plot, for both the statistical averages and derivative estimation results, a batch of $1000$ samples was directly sampled by the model at each temperature point. One can see that the statistical averages have a larger variance.
F.3 Scaling test on larger lattice sizes
Figure S6: The estimated free energy and its standard deviation (inset) of the 2D XY model on a square lattice with periodic boundary conditions. The corresponding lattice sizes are labeled in the bottom-right corner of each subfigure. The temperature factor (T-T) is removed from the free energy($T\log\mathcal{Z}$) for better comparison.
Figure S7: The mean energy and heat capacity of the 2D XY model on a square lattice with periodic boundary conditions, estimated using the differentiation of the free energy. The corresponding lattice sizes are labeled above each subfigure.
To better understand the performance of the VaTD model used in the XY model experiment, we performed a scaling test on $9$ other XY models with larger lattice sizes: $24\times 24$, $32\times 32$, $64\times 64$, $96\times 96$, $128\times 128$, $192\times 192$, $216\times 216$, $256\times 256$, and $258\times 258$. In this scaling test, the same VaTD model used in the $16\times 16$ XY model experiment in the main text is used, i.e., with the same hyperparameters and the same structure as shown in Tab. S3. Due to the GPU memory limitation (2424GB on RTX 4090), for lattice sizes larger than $32\times 32$, we used checkpointing and reduced batch size.
In Fig. S6 and Fig. S7, we present plots of the free energy, mean energy, and specific heat capacity ($C_{v}$) for the 2D XY model at various lattice sizes. These results demonstrate that the proposed VaTD training method can still identify phase transition phenomena in the 2D XY model up to a lattice size of $258\times 258$. We will need larger GPU memory or more GPU cards for a larger lattice size.
We acknowledge that a fixed neural network model tends to exhibit diminished performance with larger input sizes. However, it is important to note that the proposed VaTD method serves as a general training framework that can integrate temperature dependence into any tractable density generative model. For those targeting larger lattice sizes, we recommend using a more capable generative mode, such as the NeuralODE model [neuralode] or the continuous normalizing flow (CNF) models. Alternatively, generative models incorporating a renormalization group mechanism, as discussed in [prl, prr]. Both types of generative models can be effectively adopted within the VaTD training framework.